OpenStudy (anonymous):

I need your help! A charge of 5 x10-9 C is attracted by a charge of –3x10-7 C with a force of 0.135 N. How far apart are they? (Hint: manipulate the equation to get an expression for d)

OpenStudy (anonymous):

Coulombs law \[F = (Q_1\times Q_2)/(4\pi \times \epsilon_0 \times d^2) \]

OpenStudy (anonymous):

Could you help me on how to start it?

OpenStudy (shane_b):

You just plug in the values you know into the formula and solve for d

OpenStudy (shane_b):

\[\large 0.135N = \frac{5x10^{-9}C\times -3x10^{-7}C}{4(3.14159) \times 8.85x10^{-12} \frac{C^2}{Nm^2}\times d^2}\]As the question states, now you must manipulate this equation to solve for d.

OpenStudy (anonymous):

Thank you! I think I got it.

OpenStudy (shane_b):

np :)